26. Use algebraic properties to determine the value of the limit: VX-3-2 lim X-7
Answer #1

We have given to find the limits so when we take x=7

We get it as 0/0

So by using L hospitality rule so

lim_{x \to 7} \frac{\sqrt{x-3}-2}{x-7}

lim_{x \to 7} \frac{\frac{1}{2\sqrt{x-3}}}{1}

\frac{\frac{1}{2\sqrt{7-3}}}{1}

{\frac{1}{4}}